OpenStudy (anonymous):

What is the missing exponent? 8^-5/(8^-7)^2=8[ ]

OpenStudy (anonymous):

@Nnesha help again please?

Nnesha (nnesha):

you need to know 2 exponent rules \[\huge\rm (x^m)^n=x^{m \times n}\] and the 2nd one \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]when we divide same bases we should subtract their exponents

OpenStudy (anonymous):

so for the first one it's -7 times 2 the second one is -5 times -7? @nnesha

Nnesha (nnesha):

first one is right 2nd one you should `subtract` their exponent not multiply them

Nnesha (nnesha):

\[\frac{ 8^{-5} }{ 8^{-7 \times 2} }\] -7 times 2 = ?

OpenStudy (anonymous):

-14?

Nnesha (nnesha):

yes right so \[\frac{ 8^{-5} }{ 8^{-14}}\]

Nnesha (nnesha):

now apply 2nd exponent rule

OpenStudy (anonymous):

is it -11? I might be wrong

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ 2^3 }{ 2^{-3} } =2^{3\color{ReD}{-}(-3)}\]

Nnesha (nnesha):

we should subtract their exponents but if the exponents at the denominator is negative then you should add \[\huge\rm \frac{ 2^3 }{ 2^{-3} } =2^{3\color{ReD}{-}(-3)} =2^{3+3}\]